Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Ole Hansen
Log entry time 00:09:10 on March 24, 2011
Entry number 353018
keyword=Swing Shift Summary
Shift Crew
SL: Ole Hansen
TO: Hisham Albataineh
3rd: Werner Boeglin
Summary
Very smooth SRC 4He production running all shift, with only
minor interruptions and very high ABU.
Notable Events
- 16:00 Shift starts with production data taking in progress.
RHRS set to thetap = 33.5°,
pp = 1.3825 GeV/c. 20 cm 4He target, 4 μA
current
- 18:35 RHRS Q2 trips, apparently due to a cryo flow instability.
Resetting/cycling the magnet, which takes about 15 minutes.
- 21:45 CODA crash.
- 0:00 Shift ends with production running in progress.
Runlist
- 3009 production
- 3010 production - R-Q2 might be off for few minutes at end
- 3011 production
- 3012 production
- 3013 junk, CODA crash
- 3014 production
- 3015 production, in progress at end of shift
Total charge accumulated in runs 3009-3014: 98.55 mC